OpenStudy (lexy148):

what is the slope of the line having the equation x+1=0

OpenStudy (anonymous):

Are you sure it's x+1 = 0, or x+1 = y

OpenStudy (lexy148):

yeah im sure thats the way its written on my paper @swaschan

OpenStudy (anonymous):

Two possibilities. Either the gradient is 1, or either the gradient is 0. I say 0, because if it's x+1=0, then x=-1, and is a horizontal line intercepting -1 on the y axis, hence no gradient.

OpenStudy (lexy148):

so the slope would be -1

OpenStudy (anonymous):

-1 +1=0 when you have the same negative numbers and a positive number it always be zero. like -2+2=0, -3+3=0, 4-4=0, etc

OpenStudy (lexy148):

thx @prosper2go